OpenStudy (anonymous):

You received a $100 gift certificate to a clothing store. The store sells t-shirts for $15 and dress shirts for $23. You want to spend no more than the amount of the gift certificate. You want to leave at most $5 of the gift certificate unspent. You need at least one dress shirt. What are all of the possible combinations of t-shirts and dress shirts you could buy? a. 1 t-shirt and 4 dress shirts b. 5 t-shirts and 1 dress shirt c. 3 t-shirts and 2 dress shirts d. 2 t-shirts and 3 dress shirts

OpenStudy (anonymous):

ANSWER THIS QUESTION SO I CAN PASS MY JUNIOR YEAR!!!!!

OpenStudy (anonymous):

OMFG ANSWER THE QUESTION PLEASSSSEEEEE

OpenStudy (anonymous):

I HAVE A 23 IN THIS CLASS AND IF I DONT GET THIS RIGHT THEN I CANT PASS THE CLASS AND I FAIL MY JUNIOR YEAR.

OpenStudy (opcode):

You received a $100 gift certificate to a clothing store. The store sells t-shirts for $15 and dress shirts for $23. You want to spend no more than the amount of the gift certificate. You want to leave at most $5 of the gift certificate unspent. You need at least one dress shirt. What are all of the possible combinations of t-shirts and dress shirts you could buy? a. 1 t-shirt and 4 dress shirts b. 5 t-shirts and 1 dress shirt c. 3 t-shirts and 2 dress shirts d. 2 t-shirts and 3 dress shirts $100 (Total Money)| $15 (per t-shirt)| $23 (per dress shirt)| $5 (Needs to be left.)| Okay so: 15 * 2 = 30 15 * 3 = 45 15 * 4 = 60 15 * 5 = 75 15 * 6 = 90 (Rates of t-shirts) 23 * 1 = 23 23 * 2 = 46 23 * 3 = 69 23 * 4 = 92 23 * 5 = 115 23 * 6 = 138 (Rates of dress-shirts) 1 t-shirt and 4 dress shirts (15) + (92) = 107 (Not quite it) 5 t-shirts and 1 dress shirt (75) + 23 = 98 ( Not quite it) 3 t-shirts and 2 dress shirts (45) + (46) = 91 (Not quite it) 2 t-shirts and 3 dress shirts (30) + (69) = 99 (Not quite it) (You want to leave at most $5 of the gift certificate unspent) I'd go with "3 t-shirts and 2 dress shirts (45) + (46) = 91 (Not quite it)."